OpenStudy (anonymous):

Brown rocks cost $5 per pound. Beige rocks cost $11 per pound. How many brown rocks are needed to create a 10 lb mixture that costs $6.80 per pound?

OpenStudy (anonymous):

I think this question is much like the last one which involved the distances and their speed.

OpenStudy (anonymous):

okay so how would I set it up? x+y=z?

OpenStudy (anonymous):

I always first try to use some dimensional analysis. Have you heard of that?

OpenStudy (anonymous):

no what is it?

OpenStudy (anonymous):

In this problem our dimensions are \[\frac{ dollars }{ pound }\] But we need to know how many rocks. Is there any information that says 1 rock = x pounds? Also, knowing the dimensions allows us to pick how we want to solve for things. So if we wanted pounds, well we'd know that we would have to flip the reciprocal to get pounds up top. Does that make sense?

OpenStudy (anonymous):

yeah, so would these be my equations,x+y=10 and 5x+11y=68 and I would use substitution?

OpenStudy (anonymous):

At this point you've just created what's call a linear system of equations. You have two equations and two unknowns. You can use one to help solve the other.

OpenStudy (anonymous):

You can use any method that makes the most sense to you. Substitution is probably going to be easier here though.

OpenStudy (anonymous):

okay thanks i got it from here

OpenStudy (anonymous):

Rock on man!